OpenStudy (anonymous):

What is 36.4÷8.8 , Also what is 25.8÷8.7 , And Last what is 36.4÷8.7? I would really appreciate it if anyone would help me, THANK YOU so much if you do and god bless everyone.

OpenStudy (anonymous):

36.4/8.8= 4.13 25.8/8.7=2.96 36.4/8.7=4.18

OpenStudy (anonymous):

thank you so much , have a wonderful day (: